\begin{thebibliography}{20}

\bibitem{AKP}B. Alspach, D.L. Kreher and A. Pastine: {\it The Friedlander-Gordon-Miller Conjecture is true,} {\it Austral. J. Combin.}, {\bf 67}(2017), 11--24.

%\bibitem{AI2} B. A. Anderson and E. C. Ihrig, {\it Every finite solvable group with a unique element of order two, except the quaternion group, has a symmetric sequencing,} {\it J. Combin. Des.}, {\bf 1}(1993), 3--14.

\bibitem{JG} R. Beals, J.A. Gallian, P. Headley, and D. Jungreis, {\it Harmonious groups}, {\it J. Combin. Theory, Ser. A}, {\bf 56}(1991), no. 2, 223--238.

\bibitem{Bray} J.N. Bray, Q. Cai, P.J. Cameron, P. Spiga, and H. Zhang, {\it The Hall-Paige conjecture, and synchronization for affine and diagonal groups},  J. Algebra, {\bf 545}(2020), 27--42.

\bibitem{DJ} L. De Wolf and M. Javaheri, {\it Harmonious sequences in groups with a unique involution}, {\it J. Combin. Theory} {\bf 5}(2025), no. 3, Article 65556.

\bibitem{Evans} A. Evans, {\it The admissibility of sporadic simple groups}, J. Algebra {\bf 321}(2009), no. 1, 105--116.

\bibitem{Fr} R.J. Friedlander, B. Gordon, and M.D. Miller, {\it On a group sequencing problem of Ringel}, Congr. Numer. {\bf 21}(1978), 307--321.

\bibitem{HP} M. Hall and L. Paige, {\it Complete mappings of finite groups}, Pac. J. Math. {\bf 5}(1955), 541--549. 

\bibitem{H} P. Headly, {\it R-sequenceability and R*-sequenceability of abelian 2-groups}, Discrete Math. {\bf 131}(1994), 345--350.

%\bibitem{Jav} M. Javaheri, {\it Doubly sequenceable groups}, {\it J. Combin. Designs}, {\bf 32}(2024), no. 7, 371--387.

\bibitem{MP} A. M\"{u}esser and A. Pokrovskiy, {\it A random {H}all-{P}aige conjecture}, Invent. Math. {\bf 240}(2025), 779--867.

%\bibitem{Ollis} M. A. Ollis, {\it Sequenceable groups and related topics}, {\it Electron. J. Combin.}, DS10-May.

%\bibitem{O2} M. A. Ollis, {\it Constructing R-sequencings and terraces for groups of even order}, Algebra and Discrete Mathematics {\bf 20}(2015), no. 2, 297--316. 

\bibitem{OW} M.A. Ollis and D. T. Willmott, {\it Constructions for terraces and R-sequencings, including a proof that Bailey’s Conjecture holds for abelian groups}, J. Combin. Designs, {\bf 23}(2015), 1--17.

\bibitem{Paige} L.J. Paige, {\it Complete mappings of finite groups}, Pacific J. Math. {\bf 1}(1951), 111--116.

\bibitem{Wang} C.-D. Wang, {\it On the harmoniousness of dicyclic groups}, {\it Discrete Math.}, {\bf 120}(1993), no. 1, 221--225.

\bibitem{WL} C.-D. Wang and P.A. Leonard, {\it More on sequences in groups,} {\it Australas. J. Combin.}, {\bf 21}(2000), 187--196.

\bibitem{WL2} C.-D. Wang and P.A. Leonard, {\it On R*-sequenceability and Harmoniousness of Groups}, J. Combin. Designs, {\bf 2}(1994), no. 2, 71--78.

\bibitem{Wilcox} S. Wilcox, {\it Reduction of the Hall-Paige conjecture to sporadic simple groups}, J. Algebra, {\bf 321}(2009), no. 5, 1407--1428.

\end{thebibliography}

\documentclass{article}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{amsmath,mathabx,pifont}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\theoremstyle{remark}

\numberwithin{equation}{section}
\title{Symmetric harmoniousness of odd-order groups}
\author{ Mohammad Javaheri  \\
Department of Mathematics \\
Siena University \\
515 Loudon Road, Loudonville, NY 12211, USA\\
\small{mjavaheri@siena.edu}  
}

\date{November 2025}

\begin{document}

\maketitle

\begin{abstract}
We prove that every odd-order group is symmetric harmonious: there exists a permutation $g_0,g_1,\ldots, g_{\ell-1}$ of elements of $G$ such that the consecutive products $g_0g_1,g_1g_2,\ldots, g_{\ell-1}g_0$ also form a permutation of elements of $G$ and $g_{\ell-i}=g_i^{-1}$ for all $1\leq i \leq \ell-1$. We apply this result to obtain new examples of R*-sequenceable groups. \end{abstract}

Keywords: Harmonious groups, R*-sequenceable groups

MSC(2020): 05E16

\section{Introduction}
A {\it harmonious} sequence in a group $G$ is a permutation $g_0,g_1,\ldots, g_{\ell-1}$ of elements of $G$ such that the consecutive products $g_0g_1,g_1g_2,\ldots, g_{\ell-1}g_0$ also form a permutation of elements of $G$, where $\ell=|G|$. Such sequences were introduced by Beals et al.\ \cite{JG} in 1991 in connection with the study of complete mappings. A {\it complete mapping} of a group $G$ is a bijection $\phi: G \rightarrow G$ such that the map $x\mapsto x \phi(x)$ is also a bijection. 

If $g_0,g_1,\ldots, g_{\ell-1}$ is a harmonious sequence in $G$, then the map $\phi: g_i \mapsto g_{i+1}$ for $0\leq i \leq \ell-1$ (where $g_\ell=g_0$) is a complete mapping of $G$, and moreover $\phi$ is a single $|G|$-cycle (i.e., the orbit of every element under iteration of $\phi$ is all of $G$). Conversely, if $\phi$ is a complete mappings of $G$ that is a single $|G|$-cycle, then for every $g\in G$, the sequence $g, \phi(g), \phi^2(g),\ldots, \phi^{\ell-1}(g)$, obtained by iterating the action of $\phi$, is a harmonious sequence in $G$.

Hall and Paige \cite{HP} proved that a necessary condition for a group $G$ to admit a complete mapping is that every Sylow 2-subgroup of $G$ be either trivial or non-cyclic. They conjectured that this condition, now known as the Hall-Paige condition) is also sufficient. The Hall-Paige conjecture was affirmed in 2009 by Bray, Evans, and Wilcox \cite{Bray,Evans,Wilcox}. 

For harmonious groups, the Hall-Paige condition is not the only obstruction. Elementary 2-groups $(\mathbb{Z}_2)^n$ with $n\geq 1$ are not harmonious. Beals et al.\ \cite{JG} showed that all abelian groups except elementary 2-groups and binary groups (groups with a unique element of order 2) are harmonious \cite{JG}. 

Non-abelian harmonious groups include the dihedral and dicyclic groups of order $8n$, where $n>1$ is an integer \cite{JG,Wang,WL}. M\"{u}esser and Pokrovskiy \cite{MP} recently proved that all sufficiently large groups satisfying the Hall-Paige condition, except the elementary 2-groups, are harmonious.

If $G$ is a non-binary abelian group and $G\neq \mathbb{Z}_3$, then $G\setminus \{0\}$ is harmonious \cite{JG}. If $G$ is a binary abelian group with $\imath_G$ being the unique element of order 2, then $G\backslash \{\imath_G\}$ is harmonious \cite{DJ}. 

Wang and Leonard \cite{WL2} introduced the notion of symmetric harmonious groups and used such groups to construct new examples of R*-sequenceable groups. 

\begin{definition}
Let $G$ be a finite group.
\begin{itemize}
\item[i)] $G$ is {\it symmetric harmonious} if there exists a permutation $g_0,\ldots, g_{\ell-1}$ of elements of $G$ such that $g_0g_1,g_1g_2,\ldots, g_{\ell-1}g_0$ is also a permutation of elements of $G$ and 
$$g_ig_{\ell-i}=1_G~\mbox{for all}~1\leq i \leq \ell-1.$$
\item[ii)] $G$ is {\it R-sequenceable} if there exists a permutation $g_0,\ldots, g_{\ell-1}$ of elements of $G$ such that $g_0^{-1}g_1,g_1^{-1}g_2,\ldots, g_{\ell-1}^{-1}g_0$ is also a permutation of elements of $G$; it is R*-sequenceable if additionally $g_{0}g_{2}=g_{2}g_{0}=g_1$.
\end{itemize}
\end{definition}

Wang and Leonard \cite{WL2} proved that all odd abelian groups are symmetric harmonious. We prove the following result on symmetric harmonious groups. 

\medskip

\noindent {\bf Main Theorem.} {\it A group is symmetric harmonious if and only if it has odd order.}

\medskip 

The Hall-Paige condition is a necessary condition for the existence of R-sequencings, which yield complete mappings and orthogonal Latin squares \cite{Paige}. Friedlander et al.\ \cite{Fr} introduced R*-sequenceable groups and showed that if a Sylow 2-subgroup of $G$ is R*-sequenceable, then $G$ is R-sequenceable. 

Non-binary abelian groups with Sylow 2-subgroups of order other than 8 are R*-sequenceable \cite{Fr,H}. The groups $H$, $(\mathbb{Z}_2)^3 \times H$, and $\mathbb{Z}_2 \times \mathbb{Z}_4 \times H$, where $H$ is an odd-order abelian group satisfying certain restrictions on its Sylow 3-subgroups, are R*-sequenceable \cite{Fr,OW}. Non-abelian examples of R*-sequenceable groups include those of order the product of two distinct odd primes \cite{WL}, dihedral groups of order $4n$, where $n\geq 4$, and dicyclic groups of order $n=16,32 \pmod{48}$ \cite{WL2}.

Ollis showed that if $G$ is an even R*-sequenceable group of even order and $H$ is a nilpotent group of odd order, then $G\times H$ is R*-sequenceable. Wang and Leonard \cite[Prop.\ 1]{WL2} proved that if $G$ is an even-order R*-sequenceable group and $H$ is a symmetric harmonious group, then $G\times H$ is R*-sequenceable. Combining this with our main theorem yields an improvement of Ollis' result and provides new examples of R*-sequenceable groups.   

\begin{theorem}
If $G$ is an even R*-sequenceable group and $H$ is an odd-order group, then $G\times H$ is R*-sequenceable. 
\end{theorem}


\section{Proof of the main theorem}\label{2}

We first prove a lemma allowing us to lift symmetric harmonious sequences via normal extensions. 

\begin{lemma}\label{HinG}
Suppose $H$ is a normal subgroup of $G$ and both $H$ and $G/H$ are odd-order symmetric harmonious groups. Then $G$ is a symmetric harmonious group. 
\end{lemma}

\begin{proof}
Let $K_0,K_1,\ldots, K_{m-1}$ be a symmetric harmonious sequence in $G/H$, where each $K_i$ is a coset of $H$ in $G$ and $K_0=H$. Choose $k_i \in K_i \subseteq G$ so that $k_0=1_G$. Since $K_i K_{m-i}=H$ for $1 \leq i \leq m-1$, without loss of generality, we can choose $k_1,\ldots, k_{m-1}$ so that
\begin{equation}\label{kieq1}
k_i k_{m-i}=1_G~~\mbox{for all}~1\leq i \leq m-1.
\end{equation}
For $0\leq i \leq m-1$, we let $\mu_i$ be the product of consecutive terms $k_{i+1},\ldots, k_{m-1}$: 
$$\mu_i=k_{i+1} \cdots k_{m-1},~\mbox{for}~0\leq i \leq m-2;~\mu_{m-1}=1_G.$$
Let $t=(m-1)/2$. It follows from \eqref{kieq1} with $i=t$ that $k_tk_{t+1}=1_G$ and so 
\begin{equation}\label{kieq2}
\mu_0=k_1k_2\cdots k_tk_{t+1}\cdots k_{m-1}=k_1k_2\cdots k_{t-1}k_{t+2}\cdots k_{m-1}= \cdots=1_G.
\end{equation}
Additionally, for $1\leq r \leq m-1$, by \eqref{kieq2}, one has
\begin{align} \nonumber
k_{m-r}\mu_{m-r} & =k_{m-r}k_{m-r+1}\cdots k_{m-1}\\ \nonumber
& =k_r^{-1}k_{r-1}^{-1}\cdots k_{1}^{-1}\\ \nonumber
& =(k_1\cdots k_{r})^{-1} \\ \label{murinv}
& =k_{r+1}\cdots k_{m-1}=\mu_r.
\end{align}

Next, let $h_0,h_1,\ldots, h_{n-1}$ be a symmetric harmonious sequence in $H$. For $0\leq r \leq m-1$ and $0\leq p \leq n-1$, we define 
\begin{equation}\label{deflam}
g_{p m+r}=\begin{cases}
h_{2p}& \mbox{if $r=0$},\\
k_r  \mu_r   h_{2p+1}  \mu_r^{-1} & \mbox{if $0<r \leq m-1$},
\end{cases}
\end{equation}
where the indices in $h_{2p}$ and $h_{2p+1}$ are computed modulo $n$. Sine $H$ is a normal subgroup of $G$, we have $g_{pm+r} \in K_r$. We claim that $g_0,\ldots, g_{mn-1}$ is a symmetric harmonious sequence in $G$. 

\medskip

\noindent{\bf Step 1.} We first show that $g_0,\ldots, g_{mn-1}$ is a permutation of elements of $G$. It suffices to prove that if $i,j \in \{0,\ldots, mn-1\}$ and $g_i=g_j$, then $i=j$. Write $i=pm+r$ and $j=qm+s$ with $p,q \in \{0,\ldots, n-1\}$ and $r,s \in \{0,\ldots, m-1\}$. 

Since $g_{pm+r} \in K_r$ and $g_{qm+s}\in K_s$, the assumption $g_i=g_j$ implies that $r=s$. Suppose $r=0$. Then \eqref{deflam} gives $h_{2p}=h_{2q}$, so $2p=2q \pmod n$. As $n$ is odd, $p=q$ and $i=j$. The case of $r>0$ is similar. 

\medskip

\noindent{\bf Step 2.} Next, we need to show that if $i,j \in \{0,\ldots, mn-1\}$ and $g_ig_{i+1}=g_{j}g_{j+1}$ (where $g_{mn}=g_0$), then $i=j$. Write $i=pm+r$ and $j=qm + s$ with $p,q \in \{0,\ldots, n-1\}$ and $r,s \in \{0,\ldots, m-1\}$. Since $g_i \in K_r$, we have $g_ig_{i+1}\in K_rK_{r+1}$. 

Thus $g_ig_{i+1}=g_{j}g_{j+1}$ implies $K_{r}K_{r+1}=K_{s}K_{s+1}$ (where $K_m=K_0$). As $K_0,\ldots, K_{m-1}$ is a harmonious sequence, it follows that $r=s$. To prove $p=q$, we consider three cases:

{\it Case 1:} $r\neq 0,m-1$. Then 
\begin{align} \nonumber
g_ig_{i+1} & = k_r\mu_r h_{2p+1} \mu_r^{-1} \cdot k_{r+1} \mu_{r+1} h_{2p+1} \mu_{r+1}^{-1}\\ \nonumber
& =\mu_{r-1} (h_{2p+1})^2 \mu_{r+1}^{-1}
\end{align}
Therefore, it follows from $g_ig_{i+1}=g_jg_{j+1}$ that $h_{2p+1}^2=h_{2q+1}^2$. In an odd-order group the mapping $h\mapsto h^2$ is a bijection, hence $2p+1=2q+1 \pmod n$, and so $p=q$. 

{\it Case 2:} $r=0$. Then
\begin{align}\nonumber
g_ig_{i+1} & =h_{2p}\cdot k_1 \mu_1h_{2p+1}\mu_1^{-1},\\ \nonumber
& =(h_{2p}h_{2p+1})\mu_1^{-1}.
\end{align}
It follows from $g_ig_{i+1}=g_jg_{j+1}$ that $h_{2p}h_{2p+1}=h_{2q}h_{2q+1}$, which implies that $2p=2q \pmod n$ (since $h_0,h_1,\ldots, h_{n-1}$ is harmonious), hence $p=q$. 

{\it Case 3:} $r=m-1$. Then $i=pm+(m-1)$ and $i+1=(p+1)m +0$. Therefore,
\begin{align} \nonumber
g_ig_{i+1} & = h_{2p+1}h_{2p+2},
\end{align}
and as in Case 2, we must have $p=q$. 

\medskip

\noindent {\bf Step 3.} It is left to show that $g_ig_{mn-i}=1_G$ for all $1\leq i \leq mn$. Let $i=pm+r$ with $0\leq p \leq n-1$ and $0\leq r\leq m-1$, one has $mn-i=m(n-1-p)+m-r$. If $r>0$, then by \eqref{murinv}, one has
\begin{align} \nonumber
g_i g_{mn-i} &=k_r  \mu_r h_{2p+1} \mu_r^{-1}\cdot k_{m-r}\mu_{m-r} h_{2(n-1-p)+1} \mu_{m-r}^{-1} \\ \nonumber
& =k_r  \mu_r h_{2p+1}h_{2(n-1-p)+1} \mu_{m-r}^{-1} \\ \nonumber
& =k_r  \mu_r h_{2p+1}h_{n-(2p+1)} \mu_{m-r}^{-1} \\ \nonumber
& =k_r  \mu_r  \mu_{m-r}^{-1} \\ \nonumber
& =k_r  k_{m-r}\mu_{m-r}  \mu_{m-r}^{-1} \\ \nonumber
&=1_G.
\end{align}
If $r=0$, then $i=pm$ and $mn-i=(n-p)m$, and so
\begin{equation} \nonumber
g_i g_{mn-i} =h_{2p}  h_{2(n-p)}=h_{2p}h_{n-2p}=1_G.
\end{equation}
Thus $G$ is symmetric harmonious.  
\end{proof}

We are now ready to prove our main theorem which states that a group is symmetric harmonious if and only if it has odd order. 

\begin{proof}
We first show that if $G$ is symmetric harmonious, then it has odd order. Let $g_0,\ldots, g_{\ell-1}$ be a symmetric harmonious sequence in $G$. There exists $0\leq i \leq \ell-1$ such that $g_{i-1}g_i=1_G$ (indices modulo $\ell$), so $g_{i-1}=g_i^{-1}=g_{\ell-i}$. Thus $i-1=\ell-i \pmod \ell$, which implies that $\ell$ is odd. 

For the converse, let $G$ be an odd-order group that is not symmetric harmonious and has minimal order. Since every abelian group admits a symmetric harmonious sequence \cite{JG,WL}, $G$ is non-abelian. By Feit-Thompson theorem, $|G'|<|G|$, where $G'$ is the derived subgroup of $G$. By minimality of $|G|$, both $G'$ and $G/G'$ are symmetric harmonious, so by Lemma \ref{HinG}, $G$ is symmetric harmonious, a contradiction. Thus, every odd-order group is symmetric harmonious. 
\end{proof}

\end{document}